Perform the subnetting on the below IP Address and find out the number of networks, network of, first host, last host, and broadcast id of each network.
IP Address:
120.0.0.0/20
1
Expert's answer
2020-12-14T17:13:41-0500

Network 120.0.0.0/20 is a class B network

Since 20 bits are used for network the remaining 32-20= 12 will be used for host address.

To get the total number of hosts 2^12=4096 hosts but the first and last address are normally not used

So usable hosts will be 4096-2= 4094

To get number of networks we take total number of hosts(4096) divide by 256

4096/256= 16 networks

First usable IP address will be 120.0.0.1

Last usable IP address will be 120.0.15.254


Since broadcast address I normally the last address of the network hence: -

All 16 of the Possible /20 Networks for 120.0.x.x


Network Address        Usable Host Range                                Broadcast Address:

 

120.0.0.0            120.0.0.1 - 120.0.15.254                                     120.0.15.255

120.0.16.0           120.0.16.1 - 120.0.31.254                                  120.0.31.255

 

120.0.32.0           120.0.32.1 - 120.0.47.254                                   120.0.47.255

 

120.0.48.0           120.0.48.1 - 120.0.63.254                                    120.0.63.255

 

120.0.64.0           120.0.64.1 - 120.0.79.254                                     120.0.79.255

 

120.0.80.0           120.0.80.1 - 120.0.95.254                                      120.0.95.255

 

120.0.96.0           120.0.96.1 - 120.0.111.254                                    120.0.111.255

 

120.0.112.0        120.0.112.1 - 120.0.127.254                                   120.0.127.255

 

120.0.128.0        120.0.128.1 - 120.0.143.254                                    120.0.143.255

 

120.0.144.0        120.0.144.1 - 120.0.159.254                                    120.0.159.255

 

120.0.160.0        120.0.160.1 - 120.0.175.254                                    120.0.175.255

 

120.0.176.0        120.0.176.1 - 120.0.191.254                                   120.0.191.255

 

120.0.192.0        120.0.192.1 - 120.0.207.254                                   120.0.207.255

 

120.0.208.0        120.0.208.1 - 120.0.223.254                                   120.0.223.255

 

120.0.224.0        120.0.224.1 - 120.0.239.254                                   120.0.239.255

 

120.0.240.0        120.0.240.1 - 120.0.255.254                                   120.0.255.255